Use quotient rule to differentiate wrt to x
1/x^3 + 5

Answer by Boreal(15235)

I'm assuming this is 1/(x^3+5)
the denominator is (x^3+5)^2
the numerator is (x^3+5)*0-1(3x^2)=-3x^2
so the answer is -3x^2/(x^3+5)^2
----
can check by using product rule f(x)=1(x^3+5)^(-1)
f'(x)=-1*(x^3+5)^(-2)*3x^2, which is the answer above.